For statistics class I Have to find the probability that exactly 41 murders committed?
Simple. The answer is 0.
Over the thousands of years that humans have existed, I am certain that a lot more than 41 murders have been committed. Therefore the probability that exactly 41 murders have been committed - NOT MORE - is zero.
Simple. The answer is 0.
Over the thousands of years that humans have existed, I am certain that a lot more than 41 murders have been committed. Therefore the probability that exactly 41 murders have been committed - NOT MORE - is zero.
Simple. The answer is 0.
Over the thousands of years that humans have existed, I am certain that a lot more than 41 murders have been committed. Therefore the probability that exactly 41 murders have been committed - NOT MORE - is zero.
Simple. The answer is 0.
Over the thousands of years that humans have existed, I am certain that a lot more than 41 murders have been committed. Therefore the probability that exactly 41 murders have been committed - NOT MORE - is zero.
How many even-numbered cards in a deck of playing cards?
There are twenty even-numbered card in a standard deck of 52 playing cards, the 2, 4, 6, 8, and 10 of Spades, Diamonds, Clubs, and Hearts.
What is the probability that there are 53 Wednesdays in a leap year?
The total number of days in a leap year is 366. Then, if we want to determine the probability of 53 Wednesdays occurring in a leap year, we write 53 / 366.
What is the number of possible outcomes when a coin is tossed twice?
4 : Each coin toss multiplies the generated outcome by two. The outcome for two tosses can be Heads (H) Tails (T), H H, T T or T H. As such, with three tosses the outcome can be H H H, H H T, H T H, T H H, T T H, T T T, T H T, H T T.
We can define the number of outcomes as 2^n, where n is the amount of tosses.
How do you read a tree diagram to determine the possible outcomes for that situation?
Each leaf of the tree is one outcome. To take a simple example, consider the tree for three flips of a coin. The root splits into heads/tails, then each of the two branches splits into heads/tails, and finally each of the four branches splits into heads/tails, for 8 possible outcomes. In this simple case, each of the outcomes is equally probable. To find how many outcomes have exactly two heads, follow each branch of the tree, and put a star by the leaf each time you count two heads. You will find three leaves with exactly two heads (HHT, HTH, and THH). Add the probabilities of these three leaves, to discover that the probability of exactly three heads is 3/8.
Where can one purchase spinner bikes?
There are many companies that sell Spinner bikes. You can purchase them from Spinning themselves, or buy them from sites like Amazon or various athletic stores.
What is the probability of rolling a pair of six sided dice five times and getting four 4's?
If by 4 you mean either 3+1 or 2+2, then: (1/12)4(11/12)(5) = 55/(125) = 0.00022103
What do Independent trials in relationship to probability theory mean?
i dont know but why dont you do some research you lazy person
What is the probability of getting heads 7 times in 12 tosses of a coin?
99/512, or 19.34%.
The nCr formula can be used in this case:
12!/((12-7!)*7!) ---> 95,040/120 ---> 792
792/(2^12) = 99/512
How many different sum can you have when you roll 3 six sided dice?
Assuming they are standard dice numbers 1-6:
The minimum sum is 1 + 1 + 1 = 3
The maximum sum is 6 + 6 + = 18
So there are 18 - 3 + 1 = 16 different possible sums.
But note, that some of the sums may be obtained in more than one way, for example 4 can be obtained in 3 different ways: 1 + 1 + 2, 1 + 2 + 1, 2 + 1 + 1.
What is the probability of drawing a king or a 2 from a deck of 52 cards?
(4 + 4) in 52, or 8 in 52, or 2 in 13, or about 0.1538.
What is the probability of obtaining tails or a six?
The answer depends on the experiment: how many coins are tossed, how often, how many dice are rolled, how often.
What is the probability of getting an 8 on a roll of 2 dice?
the first roll must be anything except a 1 .. odds = 5/6
the second roll must be one specific number that adds up to 8
so its odds are = 1/6
SO . . the total odds are 5/6 * 1/6 = 5/36 = about a 14% chance.
Another way to look at it is this:
there are only 5-combinatiosn that make 8 = 2/6, 3/5, 4/4, 5/3, 6/2
and there are 36 possible combinations . . .
so, again .. the odds are 5- out of 36 rolls.
How many spades appear on the ace of spades in a standards deck of cards?
It depends on the brand, it will be either 3 or 1.
How the probability and statistics is apply in insurance?
They are used to understand how to set insurance premiums.
Take, for example, car insurance.
The probability of a new driver getting in an accident is higher than an experienced driver.
The probability of a male driver getting in an accident is higher than an female driver (in general).
The probability of a person who has had many claims filing another claim is higher than someone who has never filed a claim.
All these data guide the insurance industry to charge higher premiums to drivers who are new, male or have filed claims in the past.
What is the probability of rolling a 5 on one dice 2 times?
The probability would be that of independently rolling a five once, multiplied by the same probability for a second time.
The two rolls are treated independently with P(5) = probability of rolling a single '5' = 1/6.
P(5,5) = Probability of rolling one 5 and then another 5, out of two rolls
p(5,5)=p(5 AND 5)=p(5)*p(5)=(1/6)*(1/6)=1/36
So it's 1/36, or 0.0278
Is it unethical to purposely distort communication to get favourable outcome?
The answer depends on the situation.
In baseball, when the catcher signals the pitcher with finger signals, he is concealing his communication from the other team.
That is not unethical.
Without such justification, the answer seems to be yes.
What is the probability of guessing at least one correct answer out of ten true or fase questions?
It is 1 - (1/2)10 = 0.9990, approx.
What is the sample space when selecting a prime number between 10and30?
{11,13,17,19,23,29} are all the prime numbers between 10 and 30
When a two dice is rolled what is the probability the two number will have a sum of 8?
5/36
Out of the 36 possible and equally-probable outcomes (1st die's number and 2nd die's number), 5 combinations give you a sum equal 8:
* 1st die gives a 2 and 2nd die gives a 6
* 1st die gives a 6 and 2nd die gives a 2
* 1st die gives a 3 and 2nd die gives a 5
* 1st die gives a 5 and 2nd die gives a 3
* 1st die gives a 4 and 2nd die gives a 4
Therefore the probability that the sum is greater than 10 equals 5/36.
What does equally probable mean?
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.